# MonkeyOCR v1.5 Technical Report: Unlocking Robust Document Parsing for Complex Patterns

### v1

@@ 5.1 Traditional pipeline-based methods
Pipeline-based methods deconstruct the document parsing workflow into a sequence of specialized sub-tasks. These tasks, such as layout analysis [31; 10; 26], reading order prediction [29], text and formula detection [11; 12; 27], and table structure recognition [21; 9; 30], are each handled by an independent model. This modular design allows for flexible integration. For instance, Marker [22] supports various document formats by combining modules for OCR, layout analysis, and table recognition. It also leverages Large Language Models (LLMs) to improve cross-page table merging and inline formula parsing. Similarly, systems like MinerU [28] and ppstruct-v3 first use layout models to identify document elements. Subsequently, they employ text and formula detection models to locate the precise positions of text and formulas within these elements, which are then sent to recognition models. A final reading order prediction step assembles all recognized components into a structured output. Although these pipeline-based approaches benefit from their modularity and have achieved notable success, they are susceptible to error propagation, where inaccuracies in early stages can cascade through the system.
@@ 5.2 LMM-Based Document Parsing Models
General large multimodal models process entire document images end-to-end for content recognition, but high-resolution inputs often result in long outputs and computational inefficiency. MonkeyOCR introduces a triplet paradigm that streamlines pipeline-based methods, reduces compounding errors, and avoids the inefficiencies of full-image processing. MinerU 2.5 [18] unifies layout and reading-order prediction within a single multimodal model, while PPOCR-VL [5] adopts a similar three-stage approach, using lightweight models for structure detection and relation prediction and a large model for content recognition. OCRFlux [2] excels at page-level parsing, converting PDFs and images into clean Markdown with support for complex layouts, tables, equations, and cross-page merging. Dots.ocr [25] offers a compact multilingual solution integrating layout detection and content recognition. Nanonets-OCR2 [15] enables advanced image-to-Markdown conversion and context-aware VQA with accurate structure extraction. OlmOCR 2 [24] leverages RL-trained 7B VLMs with verifiable unit tests, achieving significant improvements in extracting tables, equations, and multi-column layouts from PDFs.
@@ 6 Conclusion
In this paper, we presented MonkeyOCR v1.5, a unified vision–language document parsing framework that advances both structural understanding and content recognition of complex documents. By jointly predicting layout and reading order through a large multimodal model and performing localized content recognition in a two-stage pipeline, MonkeyOCR v1.5 achieves a balanced trade-off between accuracy and efficiency. The proposed visual consistency–based reinforcement learning enables self-supervised refinement of table structures without the need for dense manual annotations, while the Image-Decoupled Table Parsing (IDTP) and Type-Guided Table Merging (TGTM) modules effectively resolve long-standing challenges such as embedded image handling and cross-page or cross-column table reconstruction. Experimental results on OmniDocBench v1.5 and PubTabNet, and OCRFlux-pubtabnet-single demonstrate that MonkeyOCR v1.5 surpasses state-of-the-art baselines in both accuracy and robustness. Beyond its superior performance, MonkeyOCR v1.5 is capable of handling challenging scenarios such as embedded table restoration and cross-page table merging, which are difficult for other methods. These capabilities highlight its strong potential in complex and diverse document parsing tasks.

### v2

@@ 5.1 Traditional pipeline-based methods
Pipeline-based methods deconstruct the document parsing workflow into a sequence of specialized sub-tasks. These tasks, such as layout analysis [rt-detr; layoutlmv3; yolov10], reading order prediction [layoutreader], text and formula detection [easyocr; ppocr; unimernet], and table structure recognition [tabformer; tablemaster; docgenome], are each handled by an independent model. This modular design allows for flexible integration. For instance, Marker [marker] supports various document formats by combining modules for OCR, layout analysis, and table recognition. It also leverages Large Language Models (LLMs) to improve cross-page table merging and inline formula parsing. Similarly, systems like MinerU [mineru] and ppstruct-v3 first use layout models to identify document elements. Subsequently, they employ text and formula detection models to locate the precise positions of text and formulas within these elements, which are then sent to recognition models. A final reading order prediction step assembles all recognized components into a structured output. Although these pipeline-based approaches benefit from their modularity and have achieved notable success, they are susceptible to error propagation, where inaccuracies in early stages can cascade through the system.
@@ 5.2 LMM-Based Document Parsing Models
Large Multimodal Models have significantly advanced the field of document understanding. Prior work has improved model capabilities from multiple perspectives: TextMonkey [monkey; textmonkey] enhances fine-grained visual perception through high-resolution cropping; mPLUG-DocOwl2 [mplug-docowl2] introduces cross-page modeling to enable structural reasoning across multi-page documents; and InternVL3 [internvl] leverages joint pretraining to strengthen cross-modal alignment and long-context understanding. More recently, general-purpose LMMs [qwen2.5-vl; gemini25; internvl] have been applied to document parsing by processing entire document images in a single pass. MonkeyOCR introduces a triplet paradigm that streamlines pipeline-based methods, reduces compounding errors, and avoids the inefficiencies of full-image processing. MinerU 2.5 [mineru2.5] unifies layout and reading-order prediction within a single multimodal model, while PPOCR-VL [paddleocrvl] adopts a similar three-stage approach, using lightweight models for structure detection and relation prediction and a large model for content recognition. OCRFlux [OCRFLUX] excels at page-level parsing, converting PDFs and images into clean Markdown with support for complex layouts, tables, equations, and cross-page merging. Dots.ocr [dotsocr] offers a compact multilingual solution integrating layout detection and content recognition. Nanonets-OCR2 [Nanonets-OCR2] enables advanced image-to-Markdown conversion and context-aware VQA with accurate structure extraction. OlmOCR 2 [olmocr2] leverages RL-trained 7B VLMs with verifiable unit tests, achieving significant improvements in extracting tables, equations, and multi-column layouts from PDFs.
@@ 6 Conclusion
In this paper, we presented MonkeyOCR v1.5, a unified vision–language document parsing framework that advances both structural understanding and content recognition of complex documents. By jointly predicting layout and reading order through a large multimodal model and performing localized content recognition in a two-stage pipeline, MonkeyOCR v1.5 achieves a balanced trade-off between accuracy and efficiency. The proposed visual consistency–based reinforcement learning enables self-supervised refinement of table structures without the need for dense manual annotations, while the Image-Decoupled Table Parsing (IDTP) and Type-Guided Table Merging (TGTM) modules effectively resolve long-standing challenges such as embedded image handling and cross-page or cross-column table reconstruction. Experimental results on OmniDocBench v1.5 and PubTabNet, and OCRFlux-pubtabnet-single demonstrate that MonkeyOCR v1.5 surpasses previous state-of-the-art methods in both accuracy and robustness. Beyond its superior performance, MonkeyOCR v1.5 is capable of handling challenging scenarios such as embedded table restoration and cross-page table merging, which are difficult for other methods. These capabilities highlight its potential as a foundation model for document understanding in OCRBench [ocrbench; ocrbenchv2].